Given that X N(0,1) then show that the pdf of Y=X2 is "\chi" 12

1
Expert's answer
2021-10-14T11:58:05-0400

Given that random variable "Y"is defined as "Y = X^2,"  where "X\sim N(0, 1)."

Then for "y<0," "P(Y<y)=0" and

for "y\geq 0,"

"P(Y<y)=P(X^2<y)=P(-\sqrt{y}<X<\sqrt{y})"

"=F_X(\sqrt{y})-F_X(-\sqrt{x})=F_X(\sqrt{y})-(1-F_X(\sqrt{y}))"

"=2F_X(\sqrt{y})-1"

"f_Y(y)=\dfrac{d}{dy}(2F_X(\sqrt{y})-1)=2\dfrac{dF_X(\sqrt{y})}{dy}"

"=2\dfrac{d}{dy}\bigg(\displaystyle\int_{-\infin}^{\sqrt{y}}\dfrac{1}{\sqrt{2\pi}}e^{-t^2/2}dt\bigg)"


"=2(\dfrac{1}{\sqrt{2\pi}})e^{-(\sqrt{y})^2/2}(\dfrac{1}{2\sqrt{y}})"

"=\dfrac{1}{\sqrt{2}}(\dfrac{1}{\sqrt{\pi}})e^{-y/2}(\dfrac{1}{\sqrt{y}})"

Use "\sqrt{\pi}=\Gamma(\dfrac{1}{2})." Then


"f_Y(y)=\dfrac{1}{\sqrt{2}\cdot\Gamma(\dfrac{1}{2})}y^{-1/2}e^{-y/2}"

where "F" and "f" are the cdf and pdf of the corresponding random variables.


Then "Y=X^2\sim\chi_1^2" .